5g of a gas at 27o C occupied same volume as 0.25 g of dihydrogen at 17C at same pressure. Find out molecular weight of gass.  
 

Grade:11

1 Answers

Khimraj
5 years ago
Given, moles = mass/mol.wt
➡Mass of gas = 5g, mass of hydrogen = 0.25g
➡T1­ = 300K, T2 =  290K
➡Moles of H2 = n1 = Mass/M. wt. = 0.25/2 = 0.125
➡Moles of gas = n2 = Mass/M. wt. = 5/M
➡For hydrogen P1V1 = n1RT1 …………..(i)
➡For gas P2V2 = n2RT2 …………….(ii)
➡(∵ Pressure and volume of gas |are same)
➡∴ From equation (i) and equation (ii)
➡P1V1/P1V1 = n1RT1/n2RT2 or 1 = 0.125 *300/n2 *290
➡or n2 = 0.125 *300/290 or 5/M = 0.125 *300/290
➡or 5/M = 0.129
➡∴ M = 5/0.129 = 38.76
Hope it clears.................